A certain optical fibre has attenuation of 3.5db/km at 850 nm. if 0.5 nW of optical fibre is launched into the fibre, what is the power level in mW after 4 km ?
1
Expert's answer
2019-08-16T09:19:40-0400

Attenuation is given by formula

"α=\frac {10} {L}×\log{\frac {P_{in}}{P_{out}}} (1)"

In our case, α=3.5db/km, Pin=0.5 nW, L=4 km


We have

"\u03b1\u00d7L=14 db (2)"


Using (1) and (2) we get

Pout=19.9µW


Answer

0.01991 mW